Replace the three forces shown by an equivalent force-couple system at point A. If the forces are replaced by a single resultant force, determine the distance d below point A to its line of action.
1
Expert's answer
2020-09-06T17:15:58-0400


(a) "300\cdot1.44-270\cdot0.72-210\cdot0.72=86.4(N\cdot m)"


The equivalent moment at point A "86.4(N\cdot m)"


***

For force-couple


"F\cdot d=C\cdot l\to C=F/2=360/2=180(N)" and


"l=2d=2\cdot0.24=0.48(m)"

***


(b) The equivalent force "300+270-210=360(N)"


(c) The distance "86.4\/360=0.24(m)" below point A.